# Rethinking the Text-Vision Reasoning Imbalance in MLLMs through the Lens of Training Recipes

### v1

@@ Abstract
Multimodal large language models (MLLMs) have demonstrated strong capabilities on vision-and-language tasks. However, recent findings reveal an imbalance in their reasoning capabilities across visual and textual modalities. Specifically, current MLLMs often over-rely on textual cues while under-attending to visual content, resulting in suboptimal performance on tasks that require genuine visual reasoning. We refer to this phenomenon as the modality gap, defined as the performance disparity between text-centric and vision-centric inputs. In this paper, we analyze the modality gap through the lens of training recipes. We first show that existing training recipes tend to amplify this gap. Then, we systematically explore strategies to bridge it from two complementary perspectives: data and loss design. Our findings provide insights into developing training recipes that mitigate the modality gap and promote more balanced multimodal reasoning. Our code is publicly available at https://github.com/UCSB-NLP-Chang/Bridging-Modality-Gap.
@@
Guanyu Yao1thanks:   Equal contribution., Qiucheng Wu [^1] footnotemark: 1, Yang Zhang2, Zhaowen Wang3, Handong Zhao3, Shiyu Chang1 1UC Santa Barbara  2MIT-IBM Watson AI Lab  3Adobe Research {gyao,qiucheng,chang87}@ucsb.eduyang.zhang2@ibm.com{zhawang,hazhao}@adobe.com
@@ 1 Introduction
To understand the origin of this imbalance, we focus on the training recipes of current MLLMs. An important observation is that many training samples contain overlapping information across the textual and visual modalities. In such cases, it may be easier for MLLMs to rely on the already complete textual information rather than engage in visual reasoning. We hypothesize that this training process largely contributes to the modality gap. Our preliminary evidence supports this view: under standard training setups, the gap between text and vision centric performance widens over time, underscoring the need for more balanced training strategies.
@@ 2.1 Multimodal Large Language Models
Multimodal Large Language Models (MLLMs) have emerged as powerful tools that integrate visual and textual information to perform a wide range of tasks, including image captioning, visual question answering, and geometric reasoning. Notable MLLMs include Qwen2.5-VL series Bai et al. (2025); Wang et al. (2024); Bai et al. (2023), VL-Rethinker series Wang et al. (2025a), MiniCPM series Yao et al. (2024), InternVL3.5 series Wang et al. (2025b), Kimi-VL series Team et al. (2025b), Gemma seires Team et al. (2025a), GPT5 OpenAI (2025) and Gemini Comanici et al. (2025). These models typically employ a combination of pre-trained vision encoders and large language models, fine-tuned on multimodal datasets to enhance their understanding and generation capabilities across both modalities.
@@ 2.2 Visual Reasoning in MLLMs
Qwen2.5-VL 3B PGPS9K 0.2397 0.1812 0.2105 0.0585
MathVerse 0.3568 0.2866 0.3147 0.0702
Qwen2.5-VL 7B PGPS9K 0.3775 0.2998 0.3387 0.0777
MathVerse 0.5501 0.4576 0.5110 0.0925
MiniCPM-V-4 PGPS9K 0.3470 0.3020 0.3245 0.0450
MathVerse 0.4435 0.3721 0.4007 0.0714
Gemma-3-4b-it PGPS9K 0.4050 0.2612 0.3559 0.1438
MathVerse 0.4236 0.3350 0.3705 0.0886
Kimi-VL-A3B PGPS9K 0.4057 0.3132 0.3595 0.0925
MathVerse 0.5791 0.4827 0.5123 0.0964
VL-Rethinker 7B PGPS9K 0.4045 0.3605 0.3825 0.0440
MathVerse 0.6542 0.5728 0.6053 0.0814
InternVL3.5 8B PGPS9K 0.5218 0.3965 0.4592 0.1253
MathVerse 0.6668 0.5419 0.5919 0.1249
Qwen3-VL PGPS9K 0.6970 0.6699 0.6835 0.0271
MathVerse 0.6406 0.6089 0.6167 0.0317
GPT-5222Due to API costs, the results are evaluated on a subset of 50 test samples. PGPS9K 0.9400 0.8000 0.8700 0.1400
MathVerse 0.7667 0.6333 0.7000 0.1334
Gemini 2.5 Flash [^2] Due to API costs, the results are evaluated on a subset of 50 test samples. PGPS9K 0.9200 0.7400 0.8300 0.1800
MathVerse 0.8696 0.7778 0.8200 0.0918
@@ 3 Modality Gap in MLLMs
We report the text-centric and vision-centric performance measuring on $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$. in addition, we report the overall performance as the average accuracy across $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$.
@@
Qwen2.5-VL 3B 0.2397 0.1812 0.2105
Qwen2.5-VL 3B $\mathcal{D}_{1}$ 0.6208 0.4480 0.5249
Qwen2.5-VL 3B $\mathcal{D}_{2}$ 0.5118 0.5050 0.5084
Qwen2.5-VL 7B 0.3775 0.2998 0.3387
Qwen2.5-VL 7B $\mathcal{D}_{1}$ 0.7422 0.5377 0.6400
Qwen2.5-VL 7B $\mathcal{D}_{2}$ 0.6342 0.6040 0.6191
@@ 4.1 Data Perspective
In this section, we compare the mixed training and curriculum training strategies. For mixed training settings, we stop training until the DAPO training parameter num_gen_batches reaches 10, for curriculum settings, we use the same total training steps but split evenly between the two stages, in which Stage 1 trains on $\mathcal{D}_{1}$ and Stage 2 trains on $\mathcal{D}_{2}$. More training details can be found in Appendix D.
@@
   Mixed training 0.6162 0.5683 0.5923 0.4724 0.4202 0.4411
   Curriculum Stage 1 ($\mathcal{D}_{1}$) 0.5840 0.4163 0.5002 0.4937 0.4317 0.4565
   Curriculum Stage 2 ($\mathcal{D}_{1}\!\rightarrow\!\mathcal{D}_{2}$) 0.5978 0.5400 0.5689 0.4902 0.4369 0.4582
@@
   Mixed training 0.7010 0.6565 0.6788 0.5493 0.4833 0.5097
   Curriculum Stage 1 ($\mathcal{D}_{1}$) 0.7305 0.5272 0.6289 0.4995 0.4442 0.4663
   Curriculum Stage 2 ($\mathcal{D}_{1}\!\rightarrow\!\mathcal{D}_{2}$) 0.7060 0.6630 0.6845 0.5695 0.5060 0.5314
@@ 4.2 Loss Perspective
We introduce a contrastive self-distillation KL loss to transfer reasoning ability from inputs with full text condition to those with partial text condition. For each paired prompt that shares the same image and question, $(x^{(1)},x^{(2)})\in(\mathcal{D}_{1},\mathcal{D}_{2})$, we first sample a sequence $\hat{\mathbf{y}}$ from $\pi_{\theta}(\cdot\mid x^{(1)})$. When this sequence is verified correct, we align next-token distributions under the same prefix $\hat{y}_{<t}$ by defining $p_{t}:=\pi_{\theta}(\cdot\mid\hat{y}_{<t},x^{(2)})$ and $q_{t}:=\operatorname{stopgrad}[\pi_{\theta}(\cdot\mid\hat{y}_{<t},x^{(1)})]$. We then minimize a forward KL averaged over time:
@@
We compare three training strategies in Table 4: ① Plain RL on $\mathcal{D}_{1}$, ② with KL, i.e., adding the contrastive KL loss, and ③ with KL + Curriculum, where the KL-trained model is subsequently fine-tuned on $\mathcal{D}_{2}$. From the in-distribution results on PGPS9K, both KL and KL + Curriculum consistently outperform the plain baseline, confirming that the KL term effectively transfers reasoning ability and stabilizes the optimization process. However, on the out-of-distribution dataset MathVerse, the improvements are less consistent, likely due to annotation and representation mismatches between the datasets. Specifically, PGPS9K provides explicit geometric cues, whereas MathVerse often omits such markings, weakening cross-domain transfer. We analyze this mismatch further in Appendix E. Overall, the KL loss enhances general reasoning ability, while the subsequent curriculum fine-tuning slightly degrades out-of-distribution performance, reflecting the impact of differing annotation styles across datasets.
@@ 5 Conclusion
The main limitation of this work is that our training experiments are conducted on a domain-specific dataset PGPS9K, which may limit the generalization of our findings to other datasets or tasks. Future work could explore methods to prepare both vision-centric and text-centric data based on more diverse datasets, and evaluate the effectiveness of our proposed methods on a wider range of MLLMs and tasks.
Acknowledgments
The work of Guanyu Yao, Qiucheng Wu, and Shiyu Chang was partially supported by the National Science Foundation (NSF) Grant IIS-2338252, NSF Grant IIS-2207052, and NSF Grant IIS-2302730.
@@ Appendix A Dataset Details
We adopt the open-source subset (testmini) of the MathVerse dataset as our out-of-distribution evaluation benchmark.
@@ Appendix C Complete Prompt
Appendix D Training Setup
@@
Appendix E Annotation Difference in Two Dataset

### v2

@@ Abstract
Multimodal large language models (MLLMs) have demonstrated strong capabilities on vision-and-language tasks. However, recent findings reveal an imbalance in their reasoning capabilities across visual and textual modalities. Specifically, current MLLMs often over-rely on textual cues while under-attending to visual content, resulting in suboptimal performance on tasks that require genuine visual reasoning. We refer to this phenomenon as the modality gap, defined as the performance disparity between text-centric and vision-centric inputs. In this paper, we analyze the modality gap through the lens of training recipes. We first show that existing training recipes tend to amplify this gap. Then, we systematically explore strategies to bridge it from two complementary perspectives: data and loss design. Our findings provide insights into developing training recipes that mitigate the modality gap and promote a more balanced multimodal reasoning.
@@
Guanyu Yao1thanks:   Equal contribution., Qiucheng Wu1111Due to API costs, the results are evaluated on a subset of 50 test samples., Yang Zhang2, Zhaowen Wang3, Handong Zhao3, Shiyu Chang1 1UC Santa Barbara  2MIT-IBM Watson AI Lab  3Adobe Research {gyao,qiucheng,chang87}@ucsb.eduyang.zhang2@ibm.com{zhawang,hazhao}@adobe.com
@@ 1 Introduction
To understand the origin of this imbalance, we focus on the training recipes of current MLLMs. An important observation is that many training samples contain overlapping information across the textual and visual modalities. In such cases, it may be easier for MLLMs to rely on the already complete textual information rather than engage in visual reasoning. We hypothesize that this training process largely contributes to the modality gap. Our preliminary evidence supports this view: under standard training setups, the gap between text- and vision-centric performance widens over time, underscoring the need for more balanced training strategies.
@@ 2.1 Multimodal Large Language Models
Multimodal Large Language Models (MLLMs) have emerged as powerful tools that integrate visual and textual information to perform a wide range of tasks, including image captioning, visual question answering, and geometric reasoning. Notable MLLMs include Qwen2.5-VL series Bai et al. (2025); Wang et al. (2024); Bai et al. (2023), VL-Rethinker series Wang et al. (2025a), MiniCPM series Yao et al. (2024), InternVL3.5 series Wang et al. (2025b), Kimi-VL series Team et al. (2025b), Gemma series Team et al. (2025a), GPT-5 OpenAI (2025) and Gemini Comanici et al. (2025). These models typically employ a combination of pre-trained vision encoders and large language models, fine-tuned on multimodal datasets to enhance their understanding and generation capabilities across both modalities.
@@ 2.2 Visual Reasoning in MLLMs
Qwen2.5-VL 3B PGPS9K 23.97 18.12 21.05 5.85
MathVerse 35.68 28.66 31.47 7.02
Qwen2.5-VL 7B PGPS9K 37.75 29.98 33.87 7.77
MathVerse 55.01 45.76 51.10 9.25
MiniCPM-V-4 PGPS9K 34.70 30.20 32.45 4.50
MathVerse 44.35 37.21 40.07 7.14
Gemma-3-4b-it PGPS9K 40.50 26.12 35.59 14.38
MathVerse 42.36 33.50 37.05 8.86
Kimi-VL-A3B PGPS9K 40.57 31.32 35.95 9.25
MathVerse 57.91 48.27 51.23 9.64
VL-Rethinker 7B PGPS9K 40.45 36.05 38.25 4.40
MathVerse 65.42 57.28 60.53 8.14
InternVL3.5 8B PGPS9K 52.18 39.65 45.92 12.53
MathVerse 66.68 54.19 59.19 12.49
Qwen3-VL PGPS9K 69.70 66.99 68.35 2.71
MathVerse 64.06 60.89 61.67 3.17
GPT-5111Due to API costs, the results are evaluated on a subset of 50 test samples. PGPS9K 94.00 80.00 87.00 14.00
MathVerse 76.67 63.33 70.00 13.34
Gemini 2.5 Flash [^1] Due to API costs, the results are evaluated on a subset of 50 test samples. PGPS9K 92.00 74.00 83.00 18.00
MathVerse 86.96 77.78 82.00 9.18
@@ 3 Modality Gap in MLLMs
We report the text-centric and vision-centric performance measured on $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$. In addition, we report the overall performance as the average accuracy across $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$.
@@
Qwen2.5-VL 3B 23.97 18.12 21.05
Qwen2.5-VL 3B $\mathcal{D}_{1}$ 62.08 44.80 52.49
Qwen2.5-VL 3B $\mathcal{D}_{2}$ 51.18 50.50 50.84
Qwen2.5-VL 7B 37.75 29.98 33.87
Qwen2.5-VL 7B $\mathcal{D}_{1}$ 74.22 53.77 64.00
Qwen2.5-VL 7B $\mathcal{D}_{2}$ 63.42 60.40 61.91
@@ 4.1 Data Perspective
We compare mixed training and curriculum training, matching the total training budget (details in Appendix E). Curriculum training splits steps evenly: Stage 1 on $\mathcal{D}_{1}$, followed by Stage 2 on $\mathcal{D}_{2}$.
@@
   Mixed training 57.93 52.80 55.37 46.07 41.11 43.09
   Curriculum Stage 1 ($\mathcal{D}_{1}$) 58.40 41.63 50.02 49.37 43.17 45.65
   Curriculum Stage 2 ($\mathcal{D}_{1}\!\rightarrow\!\mathcal{D}_{2}$) 59.78 54.00 56.89 49.02 43.69 45.82
@@
   Mixed training 70.10 65.65 67.88 54.93 48.33 50.97
   Curriculum Stage 1 ($\mathcal{D}_{1}$) 73.05 52.72 62.89 49.95 44.42 46.63
   Curriculum Stage 2 ($\mathcal{D}_{1}\!\rightarrow\!\mathcal{D}_{2}$) 70.60 66.30 68.45 56.95 50.60 53.14
@@ 4.2 Loss Perspective
We introduce a contrastive self-distillation KL loss to transfer reasoning from full-text ($\mathcal{D}_{1}$) to partial-text ($\mathcal{D}_{2}$) inputs. Given paired prompts $(x^{(1)},x^{(2)})$, and a correct response $\hat{\mathbf{y}}$ sampled from $\pi_{\theta}(\cdot\mid x^{(1)})$, we align the partial-text distribution $p_{t}:=\pi_{\theta}(\cdot\mid\hat{y}_{<t},x^{(2)})$ with the frozen full-text distribution $q_{t}:=\operatorname{stopgrad}[\pi_{\theta}(\cdot\mid\hat{y}_{<t},x^{(1)})]$ via a time-averaged forward KL:
@@
We compare three training strategies in Table 4: ① Plain RL on $\mathcal{D}_{1}$, ② with KL, i.e., adding the contrastive KL loss, and ③ with KL + Curriculum, where the KL-trained model is subsequently fine-tuned on $\mathcal{D}_{2}$. From the in-distribution results on PGPS9K, both KL and KL + Curriculum consistently outperform the plain baseline, confirming that the KL term effectively transfers reasoning ability and stabilizes the optimization process. However, on the out-of-distribution dataset MathVerse, the improvements are less consistent, likely due to annotation and representation mismatches between the datasets. Specifically, PGPS9K provides explicit geometric cues, whereas MathVerse often omits such markings, weakening cross-domain transfer. We analyze this mismatch further in Appendix H. Overall, the KL loss enhances general reasoning ability, while the subsequent curriculum fine-tuning slightly degrades out-of-distribution performance, reflecting the impact of differing annotation styles across datasets. Please refer to Appendix G for more comparisons with baseline methods and Appendix F for additional ablation studies.
@@ 5 Conclusion
Our proposed training recipe relies on constructing paired text-centric and vision-centric data, which currently leverages the detailed annotations of the PGPS9K geometry dataset. While we demonstrate that this method enhances performance across both domain-specific and general multimodal benchmarks (as shown in Appendix G), extending this data construction strategy to broader, less structured VQA tasks remains a challenge. Future work will explore automated methods to generate such training pairs for diverse domains, further validating the scalability of our approach.
@@ Appendix A Dataset Details
We adopt the open-source subset testmini of the MathVerse dataset as our out-of-distribution evaluation benchmark.
@@ Appendix C Complete Prompt
Appendix D Dynamic sAmpling Policy Optimization(DAPO) Overview
DAPOYu et al. (2025) serves as the base RL training method in our pipeline. For clarity, relative to the standard GRPO framework, DAPO introduces the following modifications:
  • Clip-Higher. DAPO decouples the lower and upper clipping ranges $(\epsilon_{\text{low}},\epsilon_{\text{high}})$ of the importance-sampling ratios and enlarges the upper bound. This mitigates entropy collapse and improves exploration by allowing greater flexibility for low-probability exploration tokens during policy updates.
  • Dynamic sampling. DAPO filters out prompts whose response groups are uniformly correct or uniformly incorrect, as these prompts do not generate meaningful gradients under GRPO. Since this filtering reduces the effective batch size, DAPO oversamples prompts that yield non-trivial advantages to maintain a sufficient number of gradient-contributing samples. This increases the density of useful training signals and improves the stability of GRPO-style updates.
  • Token-level policy gradient loss. In vanilla GRPO, all token log-probability terms within a response are averaged first, and this averaged value becomes the sample’s sole contribution to the update. Consequently, samples of different lengths receive equal weight, which dilutes the gradients of long chain-of-thought responses and overemphasizes short responses. DAPO eliminates this sample-level averaging and aggregates losses at the token level. Each token therefore contributes directly to the optimization objective, preventing long responses from being under-weighted and enabling finer-grained credit assignment. This produces more stable updates in long sequence RL and ensures that informative reasoning steps and error-inducing tokens are appropriately reflected in the gradient signal.
  • Overlong reward shaping. To address instability caused by excessively long or truncated responses, DAPO applies a length-aware penalty known as Soft Overlong Punishment. When a response exceeds a predefined maximum length, DAPO introduces a punishment interval in which the penalty increases smoothly with response length. Shorter responses receive no penalty, while severely overlong responses receive a fixed maximum penalty. This penalty is added to the rule-based correctness reward to discourage unnecessarily long outputs while preserving valid reasoning content. This mechanism reduces reward noise from truncation and prevents models from exploiting longer outputs to manipulate rewards.
Appendix E Training Setup
@@
Appendix F Ablation on RL under a Fixed Training Distribution
Model Text-centric ($\mathcal{D}_{1}$) Vision-centric ($\mathcal{D}_{2}$) Text-only Gap
Qwen3B Base Model 23.97 18.12 16.10 -2.02
Qwen3B Plain RL 240 56.50 46.45 46.82 0.37
Qwen3B Plain RL 320 62.08 44.80 48.70 3.90
Qwen7B Base Model 37.75 29.98 32.95 2.97
Qwen7B Plain RL 200 72.68 53.17 60.02 6.85
Qwen7B Plain RL 280 74.22 53.77 62.10 8.33
Table 5: Ablation results under a fixed training distribution. All models are trained exclusively on the Text-centric ($\mathcal{D}_{1}$) subset. The gap (Text-only - Vision-centric) increases consistently with additional RL steps, indicating that reinforcement learning disproportionately strengthens text-based reasoning.
To determine whether the observed modality gap is caused by RL training itself, rather than by a train–test distribution mismatch induced by splits of PGPS9K, we conduct a controlled evaluation in which the training distribution is strictly fixed. All models are trained exclusively on the Text-centric ($\mathcal{D}_{1}$) subset of PGPS9K, where both full textual descriptions and images are available. No modality-ablated subsets are used during training, and all RL updates are derived solely from this data.
After training, the same model checkpoints are evaluated under three inference settings: (1) Text-centric ($\mathcal{D}_{1}$), which includes both complete text and image inputs; (2) Vision-centric ($\mathcal{D}_{2}$), where textual descriptions are omitted, leaving only images and questions; and (3) Text-only, where images are withheld, providing only textual descriptions and questions. Since the training data distribution is identical across all models, performance differences across evaluation settings reflect changes in inference behavior rather than distribution mismatch.
Table 5 reports the results for Qwen models of different sizes and RL training steps. As RL progresses, performance improves consistently across all evaluation settings; however, the improvement is not uniform. Specifically, performance in the Text-only setting increases more rapidly than in the Vision-centric setting. For both 3B and 7B models, later-stage RL checkpoints exhibit higher accuracy in the Text-only setting compared to the Vision-centric setting, even though the models were never explicitly trained on text-only data.
This asymmetric improvement leads to a widening gap between text-only and vision-centric performance as RL steps increase. In contrast, base models exhibit a much smaller gap, sometimes performing comparably or even better in the Vision-centric setting. Because all models are trained on the same Text-centric ($\mathcal{D}_{1}$) dataset, these results indicate that reinforcement learning intrinsically biases optimization toward text-dominant reasoning, thereby amplifying the modality gap even when the training distribution is held constant.
Appendix G Comparison With Other Baseline Methods and General Benchmarks
To provide a direct comparison with existing baseline methods and assess the effectiveness of our proposed training strategy, we include two widely used baselines for multimodal reasoning. The first is an in-context learning baseline following MMICL Zhao et al. (2024), where four multimodal examples are provided at inference time. The second is Perception-Aware Policy Optimization (PAPO) Wang et al. (2025c), a reinforcement learning method designed to enhance multimodal perception during training. All models are trained exclusively on PGPS9K train set. All benchmarks discussed below are used solely for evaluation and comparison across different training strategies. All baselines are evaluated under the same inference protocol as our method.
We first compare these baselines on PGPS9K under both text and vision evaluation settings. Results are shown in Table 6. Across both 3B and 7B model scales, our KL+Curriculum strategy outperforms generalized base models, MMICL-style in-context learning, and PAPO.
Model Text-centric Vision-centric Avg
Qwen 2.5 7B Base Model 37.75 29.98 33.87
Qwen 2.5 7B MMICL 37.18 30.70 33.94
Qwen 2.5 7B PAPO 45.87 37.70 41.79
Qwen 2.5 7B KL+Curriculum (Ours) 73.42 67.87 70.65
Qwen 2.5 3B Base Model 23.97 18.12 21.05
Qwen 2.5 3B MMICL 24.07 20.10 22.14
Qwen 2.5 3B PAPO 39.87 33.67 36.77
Qwen 2.5 3B KL+Curriculum (Ours) 61.22 55.27 58.25
Table 6: Comparison with baseline methods on PGPS9K under Text-centric and Vision-centric evaluation settings.
We further evaluate the same set of models on MathVerse, which tests multimodal mathematical reasoning beyond the training distribution. Results in Table 7 show that our method consistently achieves the best average performance across both text and vision settings at different model scales, demonstrating that the gains are not limited to PGPS9K.
Model Text-centric Vision Avg
Qwen 2.5 7B Base Model 55.01 45.76 51.10
Qwen 2.5 7B MMICL 55.92 48.15 51.26
Qwen 2.5 7B PAPO 52.73 45.72 48.52
Qwen 2.5 7B KL (Ours) 57.00 48.79 52.07
Qwen 2.5 3B Base Model 35.68 28.66 31.47
Qwen 2.5 3B MMICL 43.80 36.21 39.24
Qwen 2.5 3B PAPO 47.87 40.57 43.49
Qwen 2.5 3B KL (Ours) 48.19 41.58 44.23
Table 7: Results on MathVerse under Text-centric and Vision-centric evaluation settings.
Finally, we evaluate the same set of models on two widely used general benchmarks, MATH500 Lightman et al. (2023) and MMSTAR Chen et al. (2024). As shown in Table 8, training on PGPS9K with different optimization strategies does not degrade performance on general QA and VQA benchmarks. In particular, our KL-regularized curriculum training achieves performance comparable to existing baselines across both benchmarks, indicating that geometry-focused training on PGPS9K does not adversely affect the model’s general reasoning and perception capabilities.
Model MATH500 MMSTAR
Qwen 2.5 7B Base Model 65.80 60.47
Qwen 2.5 7B PAPO 59.20 63.00
Qwen 2.5 7B MMICL 66.40 61.00
Qwen 2.5 7B KL+Curriculum (Ours) 62.20 61.87
Qwen 2.5 3B Base Model 54.60 50.13
Qwen 2.5 3B PAPO 61.00 54.40
Qwen 2.5 3B MMICL 64.60 49.93
Qwen 2.5 3B KL+Curriculum (Ours) 59.40 54.80
Table 8: Results on general benchmarks MATH500 and MMSTAR.
Appendix H Annotation Difference in Two Dataset